OpenStudy (anonymous):

compute derivatives: a) xe^sin(x) b) tan^3(x^2) thank you!!

OpenStudy (anonymous):

the first one is x * e^sin(x)

OpenStudy (amistre64):

x e^sins(x) + e^sin(x) = e^sin(x) (x+1)\[x e^{\sin(x)} + e^{\sin(x)} = e^{\sin(x)}(x+1)\]

OpenStudy (anonymous):

ty!

OpenStudy (amistre64):

\[\tan^3 (x^2) = 3(2x)\tan^2 (x^2) \sec^2 (x^2)\]

OpenStudy (amistre64):

that one was hard to keep track of :) might be wrong due to my inherent stupidity ;)